An acre of land is to be planted with trees 4 ft. apart. Determine number of trees needed.

Acre = 43560 sq. ft.
Not sure how to solve. Non-homework.

Answer by ikleyn(52879)

.

HINT : each the three goes with a square of land 4*4 = 16 square feet.


I hope this hint is enough for you to complete the solution on your own.


-------------
comment from student : A square of land 4 * 4 = 16 sq. ft. 43560 / 16 = 2722.5 trees Am I correct?
-------------


My response : I think you are correct and this estimation ~ 2722 trees is right.